summerdonovan149:

maryan drove to a friends house at an average rate of 55 miles per hour she took the samre route on the way back and her average rate was 45 miles per hour if she spent 2 hours driving how far did she drive in all?

summerdonovan149:

i need help

summerdonovan149:

@logan215

JamesTDG:

Just do 55+45

SmokeyBrown:

Maryan drove to and from her friends house, which would be the same distance both ways. The only difference would be how long she spent on the road during each trip. The first trip, at 55 mph, was faster than the second trip at 45 mph, by a ratio of 55/45 or 1.22... In other words, Maryann spent 1.22 times longer on the second trip compared to the first trip. If we take t to represent the time (in minutes) spent on the first trip, we can find the time she spent on either trip: 120 = t +1.22t; 120 = 2.22t; t = 54 So, Maryan spent 54 minutes driving to her friend's house and 66 minutes driving back. We can convert these times into hours, then multiply by the speeds to get the distance traveled. 54 minutes is 0.9 hours; she was going 55 mph , so the distance covered would be 55miles/hour*0.9hours = 49.5 miles. We can check to make sure the second trip was the same distance: 66 minutes is 1.1 hours; she was going 45 mph, so the distance covered would be 45*1.1 = 49.5 miles, as expected. So, the total distance traveled would be 49.5+49.5 = 99 miles